A body of mass 2 kg thrown vertically upwards with a velocity 40 m/s comes to instantaneous rest at a height 50 m. Is there any other force (other than gravitational force), if yes then calculate the value. The answer says 12 N.

Grade:11

1 Answers

Ranjita yadav
11 years ago

Yes , a force due to air drag will be there , to calculate this wehave to first find work done by air drag..

Acc. To work-energy theorem

Wgravity + Wair drag = change in K.E.

-Mgh +Wair drag = change in K.E.

Wair drag =change in K.E. +mgh

=[0-1/2mu2] +mgh = -1/2 x 2 x (40)2 + 2x 10 x 50

                            = -1600+1000 = -600J

now w.d. = F x s

F= -600/50 = -12N